Order: Cipro (ciprofloxacin) 400 mg in 200 mL DW IVPB.

The client weighs 175​ pounds, and is to receive 6​ mg/kg/h. Calculate the flow rate in drops per minute. The drop factor is 15​ gtt/mL.

A) 60 gtt/min

B) 50 gtt/min

C) 262 gtt/min

D) 1066 gtt/min
Answer #1

175 pounds = 80kg

Dose= 6mg/kg/hr

For 80Kg= 480mg/hr

200 ml contain 400 mg

1 ml contain 2 mg

So, 480 mg is present in 480/2 ml = 240 ml

Flow rate is 240ml/hr = 240*15 gtt/ hr =3600 gtt/hr = 60 gtt/min.